OpenStudy (anonymous):

What are the excluded values of the function? y=7/2x=48 a. 0.04 b. 24 c. –24 D. –0.04

OpenStudy (kropot72):

Please check your typing of the equation.

OpenStudy (anonymous):

OOPS I MEANT y = 7/2x + 48

OpenStudy (kropot72):

Is this it?\[y=\frac{7}{2}x+48\]

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

no its the image ^^

OpenStudy (lgbasallote):

i believe the excluded value will be the value that can make this undefined which is when 2x + 48 = 0

OpenStudy (lgbasallote):

so solve for x when 2x + 48 = 0

OpenStudy (lgbasallote):

that is your excluded value

OpenStudy (lgbasallote):

do you need more help?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

The value to exclude is the one that makes the denominator 0, so -24 would do that

OpenStudy (anonymous):

thanks